OpenStudy (anonymous):

The lenght of a rectangle is 2 more than twice its width. The area of the rectangle is 60 square units. What is the width of the rectangle? @SolomonZelman if you could help me please it would be great.

OpenStudy (anonymous):

It was a multiple choice question and the answers could be 12,17,6 or 5.

OpenStudy (bradely):

length of the rectangle, l =2+2b lb =60 (2+2b)b=60 2b+2b^2=60 b^2+b-30=0 b^2+6b-5b-30=0 b(b+6)-5(b+6)=0 (b+6)(b-5)=0 b=5 Source:www.mathskey.com

OpenStudy (anonymous):

Thank you ! :)